# SPHEREx Pre-Perihelion Mapping of $\mathrm{H_{2}O}$, $\mathrm{CO_{2}}$, and $\mathrm{CO}$ in Interstellar Object 3I/ATLAS

### v2

@@ I Introduction
SPHEREx (Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer) is a new NASA mission launched into low-Earth, Sun-sychronous orbit (altitude $\approx 675\,\mathrm{km}$) on 11-Mar-2025 UT (Doré et al., 2016, 2018; Bock et al., 2025). Conducting a 102-band near-infrared spectrophotometric survey of the entire sky over the course of 2 years, it was designed primarily to study (1) inflationary cosmology, (2) the history of galaxy formation, and (3) the abundance of astrobiologically important molecular ices in planet-forming regions. SPHEREx will observe and catalog measurements of everything along a line of sight on the sky—from local solar system objects (asteroids, comets, planets, KBOs) to main sequence stars, to distant quasars. One of these objects, 3I/ATLAS (3I, hereafter), was recently discovered on 01-Jul-2025 (Denneau et al., 2025) and verified within the next few days as only the 3rd ever macroscopic body detected hurtling through the solar system from the depths of interstellar space. Termed “ISOs”, bodies like 3I represent unique samples of exosystem material delivered close to Earth “for free”, allowing in-depth remote sensing analysis of their size, shape, composition, and temporal behavior during the few months when they are passing through the inner solar system.
@@
SPHEREx’s unique observational strengths lie in its ability to spectrophotometrically image an object at $0.75\mathrm{-}5.0\,\mathrm{\mu m}$ over many weeks and on large angular scales. SPHEREx’s spectrophotometric quality is attested to by its overall excellent 3I spectral match to the published $0.7\mathrm{-}5.3\,\mathrm{\mu m}$ JWST spectrum (Cordiner et al., 2025); where they differ, in the gas emission lines, it is always SPHEREx’s measurement that is the larger due to its much larger projected observing aperture (Fig. 2). (It is important to note that the area of the SPHEREx central pixel is 4 times the area of the entire JWST NIRSPEC IFU field of view.) SPHEREx’s temporal capabilities are demonstrated by the the 12 day long dust continuum reflectance lightcurve (Fig. 1c) and its large field image quality by $\sim 6\arcmin$ wide $\mathrm{CO}_{2}$ coma detection shown in Fig. 3.
@@
In this Letter we describe the complete set of August 2025 SPHEREx observations of 3I, and their data quality, reduction, and analysis. Because of these observations, we now understand that the extended optical morphology seen for 3I since its discovery is the product of robust $\mathrm{CO}_{2}$ gas sublimation that was undetectable in the first UVIS characterizations of the object reported in July 2025. We also find a very $\mathrm{CO}$-poor but $\mathrm{H_{2}O}$-ice rich cometary object with its water gas production just beginning to turn on, and evidence for thermal processing of the body as seen in highly thermalized inner solar system comets. We find good concordance with other recent imaging and spectroscopic measures of the object. We use these findings to argue that 3I must have been well-thermally processed by many close passages of its home star before ejection by a co-orbiting Neptune to stellar mass body, and that such thermal processing happens more rapidly than exosystems dynamically relax.
@@ II Observations
Ancillary lightcurve observations were obtained using the remote/robotic Solar system Objects Light-curve Observatory (SOLO; B. Lim et al. in prep), a Celestron LLC’s RASA11 telescope [^1] An 11-inch reflecting telescope manufactured by Celestron LLC located at Sierra Remote Observatory [^2] IAU observatory code G80, $37\arcdeg 4\arcmin 13.39\arcsec$ N, $119\arcdeg 24\arcmin 45.72\arcsec$ W, CA, USA in order to independently check the SPHEREx time domain trending. Combined with the currently installed imaging sensor at the prime focus, KL4040333Manufactured by FLI Kepler, a 4k by 4k CMOS Back Illuminated (BI) camera, and a UV-cut clear filter, the total throughput of the SOLO system as of writing is effectively sensitive to $\sim 0.4\mathrm{-}1.0\,\mathrm{\mu m}$, with a response closely comparable to the Gaia G band. The results are shown in Fig. 1c, and are consistent with the low (if any) level of variability detected by SPHEREx.
Ancillary spectral observations were used to independently check the SPHEREx spectral measurements were obtained at the NASA Infrared Telescope Facility (IRTF) on the nights of 09 and 10-Aug-2025 using the SpeX instrument in $R\sim 50$, $\lambda=0.7\mathrm{-}2.5\,\mathrm{\mu m}$ prism spectroscopy mode (Fig. 2). These measurements were vital for verifying initial SPHEREx spectrophotometric data reductions and quickly bootstrapping early photometric measures, and the two sets of mesurements were found to be closely consistent, especially the strong dropoff past $2.3\,\mathrm{\mu m}$ in the IRTF spectrum that was initially believed to be an instrumental artifact near the end of the spectral order.
@@ IV.2 Spectra
The measured fluxes of 3I are scaled to the geometry of $r_{\mathrm{obs}}=r_{\mathrm{hel}}=1\,\mathrm{au}$ assuming $F_{\lambda}\propto r_{\mathrm{hel}}^{-2}r_{\mathrm{obs}}^{-2}$ (Fig. 2a). The reflectance spectrum is derived by dividing the SPHEREx measurements by the solar flux. We can break up the spectrum into 5 main regions: the $0.7\mathrm{-}1.5\,\mathrm{\mu m}$ scattered light region; the $1.5\mathrm{-}2.3\,\mathrm{\mu m}$ mixed organics plus water absorption region; the $2.4\mathrm{-}3.3\,\mathrm{\mu m}$ water ice absorption region; the $3.2\mathrm{-}3.7\,\mathrm{\mu m}$ organics absorption region; and the $3.7\mathrm{-}5.0\,\mathrm{\mu m}$ mixed region. Superimposed on top of the continua are weak gas emission features due to $\mathrm{H_{2}O}$ at $2.7\mathrm{-}2.8\,\mathrm{\mu m}$ and $\mathrm{CO}$ at $4.6\mathrm{-}4.8\,\mathrm{\mu m}$, and a strong $\mathrm{CO}_{2}$ emission feature at $4.2\mathrm{-}4.4\,\mathrm{\mu m}$. The $\mathrm{CO}_{2}$ feature is so strong, in fact, that we believe we have detected not only normal $^{12}\mathrm{CO}_{2}$ centered in a doublet at $4.25$ and $4.27\,\mathrm{\mu m}$, but also a small side shoulder due to $^{13}\mathrm{CO}_{2}$ centered at $4.31\,\mathrm{\mu m}$. The agreement in spectral shape with the JWST spectrum of Cordiner et al. (2025) taken in one visit during the middle of our observational window is excellent; the main disagreement between the two spectral datasets is in the amplitude of the gas emission lines, with SPHEREx’s consistently being higher due to the much larger effective aperture of its measurements compared to the $3\arcsec\times 3\arcsec$ FOV of the JWST NIRSPEC IFU (approximately 1/4 the area of one SPHEREx pixel).
@@ IV.3 Gas Production Rates
$Q_{\mathrm{gas}}=2v_{\mathrm{gas}}rN_{\mathrm{gas}}$ (2)
where $N_{\mathrm{gas}}$ is the average column density of the gas within a circular aperture of projected radius $r$ on the sky, centered on the nucleus, and where $v_{\mathrm{gas}}$ is the gas emission velocity. We use $v_{\mathrm{gas}}=(0.85\,\mathrm{km/s})/\sqrt{r_{\mathrm{hel}}\text{ in au}}=0.48\,\mathrm{km/s}$ (Cochran et al., 2012) and $r=11,600$ km here. The average column density is given by (cf. e.g. Konno & Wyckoff, 1989)
$N_{\mathrm{gas}}=\frac{4r_{\mathrm{obs}}^{2}}{r^{2}}\frac{F_{\nu}\Delta\lambda}{gh\lambda}\left(\frac{r_{\mathrm{hel}}}{1\,\mathrm{au}}\right)^{2}$ (3)
@@
Given the SPHEREx $\mathrm{H_{2}O}$, $^{12}\mathrm{CO}_{2}$, $^{13}\mathrm{CO}_{2}$, and CO gas emission fluxes of 58, 8610, 70, and 59 mJy measured in 2 pixel radius effective apertures and corrected for 3I’s relative position to rh = 1 au, Delta = 1 au (Fig. 2), we derive gas production rates of $Q_{\mathrm{\mathrm{H_{2}O}}}=3.2\times 10^{26}\,\mathrm{molec/sec}\pm 20\,\mathrm{\%}$, $Q_{\mathrm{{}^{12}\mathrm{CO}_{2}}}=1.6\times 10^{27}\,\mathrm{molec/sec}\pm 10\%$, $Q_{\mathrm{{}^{13}\mathrm{CO}_{2}}}=1.3\times 10^{25}\,\mathrm{molec/sec}\pm 25\%$, and $Q_{\mathrm{\mathrm{CO}}}=1.0\times 10^{26}\,\mathrm{molec/sec}\pm 25\,\mathrm{\%}$. The rates of $\mathrm{CO}_{2}$ and $\mathrm{H_{2}O}$ emission are consistent with the activity of inbound thermally processed short-period Solar System comets at $3.2\,\mathrm{au}$ (Harrington Pinto et al., 2022), where water gas emission is “not fully on yet”, and the ratio of $Q_{\mathrm{{}^{13}\mathrm{CO}_{2}}}/Q_{\mathrm{{}^{12}\mathrm{CO}_{2}}}\sim 1/100$ is consistent with the ISM value 0f 70, although the $\mathrm{CO}$ production rate at $0.1Q_{\mathrm{CO}_{2}}$ is very low (see Discussion section). These updated Qgas values are also consistent with the preliminary values reported for the first 70 visits by Lisse et al. (2025), and by Cordiner et al. (2025) from their JWST observations of 06-Aug-2025 UT taken roughly halfway through the SPHEREx observing run.
@@ IV.4 Images
Images at different wavelength ranges probe the distribution of dust and important spectral features by molecules (Sect. IV.2). The large extension seen in $^{12}\mathrm{CO}_{2}$ versus $\mathrm{H_{2}O}$ is not surprising given the orders of magnitude higher measured flux emission from $\mathrm{CO}_{2}$ gas, the object’s location well within the $\sim 12\,\mathrm{au}$ $\mathrm{CO}_{2}$ ice line of the solar system, and the long lifetime ($\sim 6$ days) vs photolysis and SWCX compared to water ($\sim 1$ day). However, the falloff of the $\mathrm{CO}_{2}$ coma with respect to projected radial distance $\rho$ is definitely faster/steeper than $1/\rho$, implying that there is either acceleration of the outflow occurring or there is destruction of the outflowing $\mathrm{CO}_{2}$ going on.
@@ IV.5 Radial profiles
Only at wavelengths characteristic of emission from $\mathrm{H_{2}O}$ and $^{12}\mathrm{CO}_{2}$ do we see significant extension above stellar; the profile for $\mathrm{H_{2}O}$ is noisy enough that we can’t say much more about its power-law slope. By contrast, the $^{12}\mathrm{CO}_{2}$ profile slope is very well defined, and clearly falls off more quickly than a $\rho^{-1}$ coma, more like a $\rho^{-3/2}$ coma. This is indicative of either significant photolytic + charge exchange destruction of the $\mathrm{CO}_{2}$, or significant acceleration of the $\mathrm{CO}_{2}$. Given that the $\mathrm{CO}_{2}$ coma extends out to $\sim 3\arcmin=384,000\,\mathrm{km}$ at $3.2\,\mathrm{au}$, and $\mathrm{CO}_{2}$ gas molecules have a photolytic lifetime of $\sim 6\,\mathrm{days}=526,000\,\mathrm{s}$ and move at average speed of $0.8\,\mathrm{km/sec}$, it is hard to see how many of them could be getting destroyed close to the nucleus. Instead they are likely accelerating due to significant pressure in the gas releasing coma, an explanation supported by noting that the $\mathrm{CO}_{2}$ coma gas profiles of Fig. 3 also appears quite flat for the first $\sim 15\arcsec$ ($32,000\,\mathrm{km}$) or so before bending past a sublimation front to follow a $\rho^{-3/2}$ trend at larger distances.
@@ IV.6 Nucleus Size
Since the scattered light signal from 3I is unresolved spatially versus a point source by SPHEREx, we analyze its importance by taking limiting cases. Firstly, if we assume all the flux were to originate as sunlight scattered off of a nuclear surface of geometric albedo $p_{\mathrm{V}}=0.04$ at the phase angles $\alpha=16.3\mathrm{-}17.2\arcdeg$ of 3I during the observation, we find an effective spherical radius of the object of $\sim 23\,\mathrm{km}$ (Lisse et al., 2025). However, we know this assumption is patently false, as Hubble Space Telescope observations of the object have placed a firm upper limit, after fitting and tracing the coma flux back to the central PSF pixels, of no more than an $2.5\,\mathrm{km}$ radius, $p_{\mathrm{V}}=0.04$ point source residual. Further, no significant lightcurve has ever been measured for 3I, and a large asymmetric rotating nucleus, as all comets observed to date have shown, will produce significant lightcurve variations. And finally, the huge $\mathrm{CO}_{2}$ coma mapped by SPHEREx shows a flat-topped structure within 18” (3 pixels or 42,000 km) of the nucleus. The solution to all these issues is the same: the presence of a small, rotating, dark nucleus surrounded by a stable, approximately 100 times brighter coma. (Cometary comae are typically static in brightness over hours to days unless the nucleus produces a huge outburst of emission or has a bright focused region of jet-like activity on its surface). Solar system comets where the nucleus is greatly outshown by its coma are rare but not unheard of; they are the handful of small, “hyperactive” comets like 41P/Wirtanen and 103P/Hartley 2 (103P, hereafter) with so much icy dust in their comae that the coma solid material has more surface area than the nucleus and out-produces the nucleus in gas emission. These comets thus require high amounts of large, long lived icy particle emission, and further appear to be $\mathrm{CO}_{2}$ gas dominated and near the end of their actively outgassing lives ((Lisse et al., 2009); see the Discussion Section for further arguments for the connection between 3I and 103P). This was just the kind of behavior 3I was exhibiting during the early August 2025 SPHEREx observations reported here.

### v3

@@ I Introduction
SPHEREx (Spectro-Photometer for the History of the Universe, Epoch of Reionization and Ices Explorer) is a new NASA mission launched into low-Earth, Sun-synchronous orbit (altitude $\approx 675\,\mathrm{km}$) on 11-Mar-2025 UT (Doré et al., 2016, 2018; Bock et al., 2025). Conducting a 102-band near-infrared spectrophotometric survey of the entire sky over the course of 2 years, it was designed primarily to study (1) inflationary cosmology, (2) the history of galaxy formation, and (3) the abundance of astrobiologically important molecular ices in planet-forming regions. SPHEREx will observe and catalog measurements of everything along a line of sight on the sky—from local solar system objects (asteroids, comets, planets, KBOs) to main sequence stars, to distant quasars. One of these objects, 3I/ATLAS (3I, hereafter), was recently discovered on 01-Jul-2025 (Denneau et al., 2025) and verified within the next few days as only the 3rd ever macroscopic body detected hurtling through the solar system from the depths of interstellar space. Termed “ISOs”, bodies like 3I represent unique samples of exosystem material delivered close to Earth “for free”, allowing in-depth remote sensing analysis of their size, shape, composition, and temporal behavior during the few months when they are passing through the inner solar system.
@@
SPHEREx’s unique observational strengths lie in its ability to spectrophotometrically image an object at $0.75\mathrm{-}5.0\,\mathrm{\mu m}$ over many weeks and on large angular scales. SPHEREx’s spectrophotometric quality is attested to by its overall excellent 3I spectral match to the published $0.7\mathrm{-}5.3\,\mathrm{\mu m}$ JWST spectrum (Cordiner et al., 2025); where they differ, in the gas emission lines, it is always SPHEREx’s measurement that is the larger due to its much larger projected observing aperture (Fig. 2). (It is important to note that the area of the SPHEREx central pixel is 4 times the area of the entire JWST NIRSPEC IFU field of view.) SPHEREx’s temporal capabilities are demonstrated by the 12 day long dust continuum reflectance lightcurve (Fig. 1c) and its large field image quality by $\sim 6\arcmin$ wide $\mathrm{CO}_{2}$ coma detection shown in Fig. 3.
@@
In this Letter we describe the complete set of August 2025 SPHEREx observations of 3I, and their data quality, reduction, and analysis. Because of these observations, we now understand that the extended optical morphology seen for 3I since its discovery is the product of robust $\mathrm{CO}_{2}$ gas sublimation that was undetectable in the first UVIS characterizations of the object reported in July 2025. We also find a very $\mathrm{CO}$-poor but $\mathrm{H_{2}O}$-ice rich cometary object with its water gas production just beginning to turn on, and evidence for thermal processing of the body as seen in highly thermalized inner solar system comets. We find good concordance with other August 2025 imaging and spectroscopic measures of the object. We use these findings to argue that the surface layers of 3I must have been well-thermally processed before its ejection from its home system into the ISM, and that such thermal processing happens more rapidly than exosystems dynamically relax.
@@ II Observations
Ancillary lightcurve observations were obtained using the remote/robotic Solar system Objects Light-curve Observatory (SOLO; B. Lim et al., submitted to JKAS), a Celestron LLC’s RASA11 telescope [^1] An 11-inch reflecting telescope manufactured by Celestron LLC located at Sierra Remote Observatory [^2] IAU observatory code G80, $37\arcdeg 4\arcmin 13.39\arcsec$ N, $119\arcdeg 24\arcmin 45.72\arcsec$ W, CA, USA in order to independently check the SPHEREx time domain trending. Combined with the currently installed imaging sensor at the prime focus, KL4040333Manufactured by FLI Kepler, a 4k by 4k CMOS Back Illuminated (BI) camera, and a UV-cut clear filter, the total throughput of the SOLO system as of writing is effectively sensitive to $\sim 0.4\mathrm{-}1.0\,\mathrm{\mu m}$, with a response closely comparable to the Gaia G band. The results are shown in Fig. 1c, and are consistent with the low (if any) level of variability detected by SPHEREx.
Ancillary spectral observations were used to independently check the SPHEREx spectral measurements were obtained at the NASA Infrared Telescope Facility (IRTF) on the nights of 09 and 10-Aug-2025 using the SpeX instrument in $R\sim 50$, $\lambda=0.7\mathrm{-}2.5\,\mathrm{\mu m}$ prism spectroscopy mode (Fig. 2). These measurements were vital for verifying initial SPHEREx spectrophotometric data reductions and quickly bootstrapping early photometric measures, and the two sets of measurements were found to be closely consistent, especially the strong dropoff past $2.3\,\mathrm{\mu m}$ in the IRTF spectrum that was initially believed to be an instrumental artifact near the end of the spectral order.
@@ IV.2 Spectra
The measured fluxes of 3I are scaled to the geometry of $r_{\mathrm{obs}}=r_{\mathrm{hel}}=1\,\mathrm{au}$ assuming $F_{\lambda}\propto r_{\mathrm{hel}}^{-2}r_{\mathrm{obs}}^{-2}$ (Fig. 2a). The reflectance spectrum is derived by dividing the SPHEREx measurements by the solar flux. We can break up the spectrum into 5 main regions: the $0.7\mathrm{-}1.5\,\mathrm{\mu m}$ scattered light region; the $1.5\mathrm{-}2.3\,\mathrm{\mu m}$ mixed organics plus water absorption region; the $2.4\mathrm{-}3.3\,\mathrm{\mu m}$ water ice absorption region; the $3.2\mathrm{-}3.7\,\mathrm{\mu m}$ organics absorption region; and the $3.7\mathrm{-}5.0\,\mathrm{\mu m}$ mixed region. Superimposed on top of the continua are weak gas emission features due to $\mathrm{H_{2}O}$ at $2.7\mathrm{-}2.8\,\mathrm{\mu m}$ and $\mathrm{CO}$ at $4.6\mathrm{-}4.8\,\mathrm{\mu m}$, and a strong $\mathrm{CO}_{2}$ emission feature at $4.2\mathrm{-}4.4\,\mathrm{\mu m}$. The $\mathrm{CO}_{2}$ feature is strong enough that we believe we have detected not only normal $^{12}\mathrm{CO}_{2}$ centered in a doublet at $4.25$ and $4.27\,\mathrm{\mu m}$, but also a small side shoulder due to $^{13}\mathrm{CO}_{2}$ centered at $4.31\,\mathrm{\mu m}$. The agreement in spectral shape with the JWST spectrum of Cordiner et al. (2025) taken in one visit during the middle of our observational window is excellent; the main disagreement between the two spectral datasets is in the amplitude of the gas emission lines, with SPHEREx’s consistently being higher due to the much larger effective aperture of its measurements compared to the $3\arcsec\times 3\arcsec$ FOV of the JWST NIRSPEC IFU (approximately 1/4 the area of one SPHEREx pixel).
@@ IV.3 Gas Production Rates
$Q_{\mathrm{gas}}=2\pi v_{\mathrm{gas}}\rho N_{\mathrm{gas}}$ (2)
where $N_{\mathrm{gas}}$ is the average column density of the gas within a circular aperture of projected radius $\rho$ on the sky, centered on the nucleus, and where $v_{\mathrm{gas}}$ is the gas emission velocity. We use $v_{\mathrm{gas}}=(0.85\,\mathrm{km/s})/\sqrt{r_{\mathrm{hel}}\text{ in au}}=0.48\,\mathrm{km/s}$ (Cochran et al., 2012) and $r=11,600$ km here. The average column density is given by (cf. e.g. Konno and Wyckoff, 1989)
$N_{\mathrm{gas}}=\frac{4r_{\mathrm{obs}}^{2}}{\rho^{2}}\frac{F_{\nu}\Delta\lambda}{gh\lambda}\left(\frac{r_{\mathrm{hel}}}{1\,\mathrm{au}}\right)^{2}$ (3)
@@
Given the SPHEREx $\mathrm{H_{2}O}$, $^{12}\mathrm{CO}_{2}$, $^{13}\mathrm{CO}_{2}$, and CO gas emission fluxes of 58, 8610, 70, and 59 mJy measured in 2 pixel radius effective apertures and corrected for 3I’s relative position to $r_{\mathrm{hel}}=1\,\mathrm{au}$, $r_{\mathrm{obs}}=1\,\mathrm{au}$ (Fig. 2), we derive gas production rates of $Q_{\mathrm{\mathrm{H_{2}O}}}=3.2\times 10^{26}\,\mathrm{molec/sec}\pm 20\,\mathrm{\%}$, $Q_{\mathrm{{}^{12}\mathrm{CO}_{2}}}=1.6\times 10^{27}\,\mathrm{molec/sec}\pm 10\%$, $Q_{\mathrm{{}^{13}\mathrm{CO}_{2}}}=1.3\times 10^{25}\,\mathrm{molec/sec}\pm 25\%$, and $Q_{\mathrm{\mathrm{CO}}}=1.0\times 10^{26}\,\mathrm{molec/sec}\pm 25\,\mathrm{\%}$. The rates of $\mathrm{CO}_{2}$ and $\mathrm{H_{2}O}$ emission are consistent with the activity of inbound thermally processed short-period Solar System comets at $3.2\,\mathrm{au}$ (Harrington Pinto et al., 2022), where water gas emission is “not fully on yet”, and the ratio of $Q_{\mathrm{{}^{13}\mathrm{CO}_{2}}}/Q_{\mathrm{{}^{12}\mathrm{CO}_{2}}}\sim 1/100$ is consistent with the ISM value 0f 70, although the $\mathrm{CO}$ production rate at $0.1Q_{\mathrm{CO}_{2}}$ is very low (see Discussion section). These updated Qgas values are also consistent with the preliminary values reported for the first 70 visits by Lisse et al. (2025), and by Cordiner et al. (2025) from their JWST observations of 06-Aug-2025 UT taken roughly halfway through the SPHEREx observing run.
@@ IV.4 Images
Images at different wavelength ranges probe the distribution of dust and important spectral features by molecules (Sect. IV.2). The large extension seen in $^{12}\mathrm{CO}_{2}$ versus $\mathrm{H_{2}O}$ is not surprising given the orders of magnitude higher measured flux emission from $\mathrm{CO}_{2}$ gas, the object’s location well within the $\sim 12\,\mathrm{au}$ $\mathrm{CO}_{2}$ ice line of the solar system, and the long lifetime ($\sim 6$ days) vs photolysis and SWCX compared to water ($\sim 1$ day). However, the falloff of the $\mathrm{CO}_{2}$ coma with respect to projected radial distance $\rho$ is definitely steeper than $1/\rho$, implying that there is either acceleration of the outflow occurring or there is destruction of the outflowing $\mathrm{CO}_{2}$ going on.
@@ IV.5 Radial profiles
Only at wavelengths characteristic of emission from $\mathrm{H_{2}O}$ and $^{12}\mathrm{CO}_{2}$ do we see significant extension above stellar; the profile for $\mathrm{H_{2}O}$ is noisy enough that we cannot say much more about its power-law slope. By contrast, the $^{12}\mathrm{CO}_{2}$ profile slope is very well defined, and clearly falls off more quickly than a $\rho^{-1}$ coma, more like a $\rho^{-3/2}$ coma. Assuming that the CO2 coma is optically thin, (valid for the observed QCO2 = 1 x 1027 emission rate into r ¿6”) this is indicative of either significant photolytic + charge exchange destruction of the $\mathrm{CO}_{2}$, or significant acceleration of the $\mathrm{CO}_{2}$. Given that the $\mathrm{CO}_{2}$ coma extends out to $\sim 3\arcmin=384,000\,\mathrm{km}$ at $3.2\,\mathrm{au}$, and $\mathrm{CO}_{2}$ gas molecules with a photolytic lifetime of $\sim 6\,\mathrm{days}=526,000\,\mathrm{s}$ must move at an average speed of $0.7\,\mathrm{km/sec}$, it is hard to see how many of them could be getting destroyed close to the nucleus. Instead they are likely accelerating due to significant pressure in the gas releasing coma, an explanation supported by noting that the (Cochran et al., 2012) gas velocity relation has the molecules emitted thermally at 0.48 km/sec (Section 4.3), and that the $\mathrm{CO}_{2}$ coma gas profiles of Fig. 3 appear quite flat for the first $\sim 15\arcsec$ ($32,000\,\mathrm{km}$) before bending past a sublimation front to follow a $\rho^{-3/2}$ trend at larger distances.
@@ IV.6 Nucleus Size
Since the scattered light signal from 3I is unresolved spatially versus a point source by SPHEREx, we analyze its importance by taking limiting cases. Firstly, if we assume all the flux were to originate as sunlight scattered off of a nuclear surface of geometric albedo $p_{\mathrm{V}}=0.04$ at the phase angles $\alpha=16.3\mathrm{-}17.2\arcdeg$ of 3I during the observation, we find an effective spherical radius of the object of $\sim 23\,\mathrm{km}$ (Lisse et al., 2025). However, we know this assumption is patently false, as Hubble Space Telescope observations of the object have placed a firm upper limit, after fitting and tracing the coma flux back to the central PSF pixels, of no more than an $2.5\,\mathrm{km}$ radius, $p_{\mathrm{V}}=0.04$ point source residual. Further, no significant lightcurve has ever been measured for 3I, and a large asymmetric rotating nucleus, as all comets observed to date have shown, will produce significant lightcurve variations. And finally, the giant $\mathrm{CO}_{2}$ coma mapped by SPHEREx shows a flat-topped structure within $15\arcsec$ (2.5 pixels or 32,000 km) of the nucleus.
The solution to all these issues is the same: the presence of a small, rotating, dark nucleus surrounded by a stable, approximately 100 times brighter coma. (Cometary comae are typically static in brightness over hours to days unless the nucleus produces a huge outburst of emission or has a bright focused region of jet-like activity on its surface). Solar system comets where the nucleus is greatly outshown by its coma are rare but not unheard of; they are the handful of small, “hyperactive” comets like 41P/Wirtanen and 103P/Hartley 2 (103P, hereafter) with so much icy dust in their comae that the coma solid material has more surface area than the nucleus and out-produces the nucleus in gas emission. These comets thus require high amounts of large, long lived icy particle emission, and further appear to be $\mathrm{CO}_{2}$ gas dominated ((Lisse et al., 2009); see the Discussion Section for further arguments for the connection between 3I and 103P). This was just the kind of behavior 3I was exhibiting during the early August 2025 SPHEREx observations reported here.